# The Interplay Between Logical Phenomena and the Cognitive System of the Mind

### 08 Jan 2024 v1

@@ 4. How are logical and conceptual phenomena extracted from experiential phenomena?
When a system emerges as a consequence of physical and chemical factors within its environment, the logic and regularity observed within the system are intricately connected to the underlying physical and chemical properties of that environment. In other words, the behavior and functioning of the system’s components are directly influenced by the mobility and characteristics of the surrounding environment. As the system interacts with its environment, it becomes attuned to the patterns and cues present within its surroundings, perceiving a certain degree of regularity. This perception of regularity enables the system to exhibit intelligent adaptation and responsiveness based on the available environmental cues. Through the interplay between the system and its environment, the system can discern meaningful information, recognize recurring patterns, and adjust its behaviors or internal states accordingly. This adaptation allows the system to optimize its functioning and enhance its chances of survival or achieving its objectives within the given environmental context. It is important to note that the notion of intelligent adaptation and perception of regularity does not necessarily imply conscious awareness or intentionality. Instead, it reflects the inherent interconnectedness between the system and its environment, where the system’s behavior emerges from the dynamic interactions between its components and the external factors at play. This perspective acknowledges the profound influence that the physical and chemical properties of the environment exert on the behavior and functioning of systems, highlighting the intricate relationship between a system’s internal processes and its external context.
@@
According to the previous section, we know that the dynamic kinetic stability and the semblance of life are reliant on three fundamental properties ($TFP$). The manifestation of characteristic $TFP$ in living organisms, spanning from single-celled organisms to humans, is attributed to the function of $ASAG$. The evolutionary progression of sensors has consequently influenced the development of $TFP$, thereby enhancing the capabilities of sensors within the systems we encounter. As a logical outcome, the processes occurring within the system are a consequence of the chemical and physical activities involved in the creation and evolution of $TFP$.
@@
The establishment of this logic is a common outcome across various living organisms, arising from the interplay between their internal system and the behavior of their surrounding environment. In this context, ”correspondence” refers to the chemical and physical reactions occurring within the system, which are directly influenced by the behavior of the surrounding environment. In essence, the system’s activity is a function of the behavior exhibited by its environment. Put differently, the chemical and physical changes taking place within the systems have established a correspondence with their surroundings, serving to maintain and stabilize the state of $TFP$, according to the paper of Pross (2021)[55]. Consequently, this has contributed to the emergence of a molecular, chemical, and physical logic (or quasi-logic) within the systems, driven by various natural forces. Indeed, the interaction of forces and the establishment of a unique state on Earth and in the solar system have played a role in initiating chemical and physical processes within the system, ultimately leading to the creation of $L_{TFP}$ within the system.
The development of human conscious experience is facilitated by the emergence and activation of new sensors within the system, see Azar (2023)[4]. As the process of complete system consciousness unfolds, it gives rise to pseudo-logical behavior ($L_{TFP}$), which can lead to illusions, heightened skills, and other abilities in individuals. For instance, phenomena such as love, falling in love, or experiencing pleasure from food can be attributed to the development of the $L_{TFP}$ mechanism. However, the development of new sensors can also give rise to illusions in humans and certain animals, sometimes leading to challenges in their relationship with $L_{TFP}$. Examples of these challenges include instances of humans engaging in harmful actions like killing children or self-starvation due to delusional beliefs, as well as cases of animals exhibiting self-destructive behaviors. These actions can be attributed to the activity of newly developed sensors. Throughout history, the occurrence of illusions in humans has increased, and this can be directly linked to the expansion of human consciousness potential specifically, the birth of new sensors as part of the $ASAG$ system. In this context, an illusion refers to an emotional response to an event or internal thoughts and feelings that are largely unrelated to the event itself or our true desires. For instance, when a person encounters an internal or external event (denoted as $\alpha$), they may attribute an emotional cause to it that is not the primary cause of event $\alpha$. Furthermore, they may propose a solution to address this emotional cause that is completely unrelated to the actual event.
@@
In the realm of logical reasoning, there are certain propositions that are considered axioms. These propositions are typically self-evident or universally accepted and serve as the foundation for further reasoning. They are not derived from other propositions but are treated as valid starting points. For instance, propositions that are equivalent to a valid statement such as $q\vee\sim q$ (the law of excluded middle) are considered axioms. These propositions are inherently valid based on their logical structure. On the other hand, we cannot assert with certainty that an event will or will not occur in a specific moment. Therefore, any proposition that leads to a contradiction, such as $p\wedge\sim p$ (the law of non-contradiction), is considered invalid. In summary, in order to establish the validity of a proposition like $p_{0}$, we need to examine its logical structure, determine if it is an axiom or derived from axioms, and ensure that it does not lead to a contradiction. By adhering to these principles of logical reasoning, we can evaluate the validity of propositions and construct a reliable framework for knowledge and inference.
@@
    By quality of our conscious experience.
@@ 3. $L_{3}$:
If we conceive of all living beings as belonging to a set $\Sigma$, we can establish a function $f$ that maps each living being in $\Sigma$ to the interval $[0,+\infty)$. Here, the value of $f(x)$ represents the degree of consciousness exhibited by the living being $x$. Consequently, each living being possesses a unique logic that is contingent upon its specific degree of consciousness. To illustrate this concept, let us consider the examples of a human ($x$) and a cat ($y)$. It is widely acknowledged that humans, in comparison to cats, exhibit a higher level of consciousness when confronted with various phenomena. For instance, humans possess the cognitive ability to comprehend the precise causes of an earthquake through a particular logical framework, whereas a cat lacks this capacity. Therefore, by virtue of their respective degrees of consciousness, we can infer that $f(y)<f(x)$, signifying that the degree of consciousness for a cat ($y$) is lower than that of a human ($x$). This observation highlights the varying levels of consciousness across different living beings, leading to distinct modes of logic shaped by their individual cognitive capacities.

### 21 Jan 2024 v2

@@ 4. Extracting Logical and Conceptual Phenomena from Experiential Realms
When a system arises due to the interplay of physical and chemical factors in its environment, the logic and regularity observed within the system are closely intertwined with the underlying physical and chemical properties of that environment. In simpler terms, the behavior and functioning of the system’s components are directly influenced by the mobility and characteristics of the surrounding environment, as well as the internal chemical and physical reactions of the system that interact with it. As the system interacts with its environment, it becomes attuned to the patterns and cues present within its surroundings, perceiving a certain degree of regularity. This perception of regularity enables the system to exhibit intelligent adaptation and responsiveness based on the available environmental cues. Through the interplay between the system and its environment, the system can discern meaningful information, recognize recurring patterns, and adjust its behaviors or internal states accordingly. This adaptation allows the system to optimize its functioning and enhance its chances of survival or achieving its objectives within the given environmental context. It is important to note that the notion of intelligent adaptation and perception of regularity does not necessarily imply conscious awareness or intentionality. Instead, it reflects the inherent interconnectedness between the system and its environment, where the system’s behavior emerges from the dynamic interactions between its components and the external factors at play. This perspective acknowledges the profound influence that the physical and chemical properties of the environment exert on the behavior and functioning of systems, highlighting the intricate relationship between a system’s internal processes and its external context.
@@
According to the Section 2, we know that the dynamic kinetic stability and the semblance of life are reliant on three fundamental properties ($TFP$). The manifestation of characteristic $TFP$ in living organisms, spanning from single-celled organisms to humans, is attributed to the function of $ASAG$. The evolutionary progression of sensors has consequently influenced the development of $TFP$, thereby enhancing the capabilities of sensors within the systems we encounter. As a logical outcome, the processes occurring within the system are a consequence of the chemical and physical activities involved in the creation and evolution of $TFP$.
@@
The establishment of this logic is a common outcome across various living organisms, arising from the interplay between their internal system and the behavior of their surrounding environment. In this context, ”correspondence” refers to the chemical and physical reactions occurring within the system, which are directly influenced by the behavior of the surrounding environment. In essence, the system’s activity is a function of the behavior exhibited by its environment. Put differently, the chemical and physical changes taking place within the systems have established a correspondence with their surroundings, serving to maintain and stabilize the state of $TFP$, according to the paper of Pross (2021)[60]. Consequently, this has contributed to the emergence of a molecular and chemical-physical logic (or quasi-logic) within the systems, driven by various natural forces. Indeed, the interaction of forces and the establishment of a unique state on Earth and in the solar system have played a role in initiating chemical and physical processes within the system, ultimately leading to the creation of $L_{TFP}$ within the system.
The development of human conscious experience is facilitated by the emergence and activation of new sensors within the system, see Azar (2023)[5]. As the process of complete system consciousness unfolds, it gives rise to pseudo-logical behavior ($L_{TFP}$), which can lead to illusions, heightened skills, and other abilities in individuals. For instance, phenomena such as love, falling in love, or experiencing pleasure from food can be attributed to the development of the $L_{TFP}$ mechanism. However, the development of new sensors can also give rise to illusions in humans and certain animals, sometimes leading to challenges in their relationship with $L_{TFP}$. Examples of these challenges include instances of humans engaging in harmful actions like killing children or self-starvation due to delusional beliefs, as well as cases of animals exhibiting self-destructive behaviors. These actions can be attributed to the activity of newly developed sensors. Throughout history, the occurrence of illusions in humans has increased, and this can be directly linked to the expansion of human consciousness potential specifically, the birth of new sensors as part of the $ASAG$ system. In this context, an illusion refers to an emotional response to an event or internal thoughts and feelings that are largely unrelated to the event itself or our true desires. For instance, when a person encounters an internal or external event, they may attribute an emotional cause to it that is not the primary cause of the event. Furthermore, they may propose a solution to address this emotional cause that is completely unrelated to the actual event.
@@
In the realm of logical reasoning, there are certain propositions that are considered axioms. These propositions are typically self-evident or universally accepted and serve as the foundation for further reasoning. They are not derived from other propositions but are treated as valid starting points. For instance, propositions that are equivalent to a valid statement such as $p\vee\sim p$ (the law of excluded middle) are considered axioms. These propositions are inherently valid based on their logical structure. On the other hand, we cannot assert with certainty that an event will or will not occur in a specific moment. Therefore, any proposition that leads to a contradiction, such as $p\wedge\sim p$, is considered invalid. In summary, in order to establish the validity of a proposition like $p_{0}$, we need to examine its logical structure, determine if it is an axiom or derived from axioms, and ensure that it does not lead to a contradiction. By adhering to these principles of logical reasoning, we can evaluate the validity of propositions and construct a reliable framework for knowledge and inference.
@@
    Role of Conscious Experience.
By engaging with these two fundamental influences, humans develop a logic that is informed by the forces of nature and the depth of conscious experience. This intertwined relationship between external observations and internal reflections allows for the continuous refinement and evolution of human logic.
@@ 3. $L_{3}$:
The degree of consciousness experience refers to the level or intensity of awareness and subjective experience that an individual possesses. It represents the extent to which a person is conscious of their surroundings, thoughts, emotions, and sensory perceptions. The degree of consciousness experience varies among different systems, for more information, please refer to Tononi (2004)[65], Tononi and Christof (2015)[71] and Lee (2023)[41].
If we classify living organisms into distinct categories, such as humans in one category and cats or ants in others, and denote the set of these categories as $\Sigma$, we can establish a function $f$ that maps each living being in $\Sigma$ to the interval $[0,+\infty)$. In this context, the value of $f(x)$ represents the level of consciousness exhibited by the living being $x$. This level of consciousness encompasses the cognitive abilities and mental experiences of $x$, which contribute to its capacity for perceiving reality and engaging in logical analysis of phenomena. Consequently, each living being possesses a unique logic that is contingent upon its specific degree of consciousness. To illustrate this concept, let us consider the examples of a human ($x$) and a cat ($y)$. It is widely acknowledged that humans, in comparison to cats, exhibit a higher level of consciousness when confronted with various phenomena. For instance, humans possess the cognitive ability to comprehend the precise causes of an earthquake through a particular logical framework, whereas a cat lacks this capacity. Therefore, by virtue of their respective degrees of consciousness, we can infer that $f(y)<f(x)$, signifying that the degree of consciousness for a cat ($y$) is lower than that of a human ($x$). This observation highlights the varying levels of consciousness across different living beings, leading to distinct modes of logic shaped by their individual cognitive capacities.